Camel tests and examples^{1}

We currently have support for articles (including newspaper articles), items in collections of essays, books, sections of books, ephemeral booklets, masters theses, Commonwealth, US and Japanese cases, and statutes from a few jurisdictions. An example of each is given below, first without, then with a pinpoint. This does not give a complete picture of the ‘formatting tree’, it”s just a sample.

• Ordinary articles

– Macauley, Non-Contractual Relations in Business: A Preliminary Study, 28 Am. Soc. Rev. 55 (1963)

– Macauley, Non-Contractual Relations in Business: A Preliminary Study, 28 Am. Soc. Rev. 55, 56 (1963)

• Newspaper articles

– Appleyard, Why Major should heed the press, The Independent, p. 21 Jul. 7, 1995

– Appleyard, Why Major should heed the press, The Independent, p. 21, 21 Jul. 7, 1995 (A silly example, since there’s only one page to the piece!)

• Articles in collections of essays

– Haley, Japan’s New Land and House Lease Law, in Land Issues in Japan: A Policy Failure? 149 (1992)

– Haley, Japan’s New Land and House Lease Law, in Land Issues in Japan: A Policy Failure? 149, 152 (1992)

• Books

– M. Goosens, F. Mittelbach, and A. Samarin, The L^{A}TEX Companion: The L^{A}TEX Companion (1994)

– M. Goosens, F. Mittelbach, and A. Samarin, The L^{A}TEX Companion: The L^{A}TEX Companion 371 (1994)

• Sections of books

– M. Goosens, F. Mittelbach, and A. Samarin, Bibliography Generation, Chapter 13 of The L^{A}TEX Companion 371 (1994)

– M. Goosens, F. Mittelbach, and A. Samarin, Bibliography Generation, Chapter 13 of The L^{A}TEX Companion, 371, 374–75 (1994) (The BibTeX processing flow shown here is simplified for Camel users)

1Note that, in addition to admiring the examples in this document, you can tinker with it to produce different kinds of bibliographies. See the comments in the file for suggestions.

• Ephemeral booklets

– Sir George Bailey Sansom, The First Japanese Constitution: A Lecture (1938)

– Sir George Bailey Sansom, The First Japanese Constitution: A Lec-

ture 2 (1938)

• Masters theses

– Masahiro Homma, The Derivative Suit under Japanese Law as a Measure of Corporate Governance (LL.M. thesis, School of Oriental and African Studies, 1994)

– Masahiro Homma, The Derivative Suit under Japanese Law as a Measure of Corporate Governance 23 (LL.M. thesis, School of Oriental and African Studies, 1994)

• Commonwealth law cases – Heap v Motorists’ Advisory Agency Ltd, [1923] 1 KB 577

– Heap v Motorists’ Advisory Agency Ltd, [1923] 1 KB 577, 578^{2}

• US law cases – Bradshaw v United States, 231 Ct.Cl. 144; 683 F.2d 365 (1982)

– Bradshaw v United States, 231 Ct.Cl. 144, 145; 683 F.2d 365, 366

(1982)

Titles that end in a numeral are a special treat. Camel is not tricked by them.

• J. M. Maki, Court and Constitution in Japan: Selected Supreme Court Decisions, 1948-1960, 23 (1964)

• J. M. Maki, Court and Constitution in Japan: Selected Supreme

Court Decisions, 1948-1960 (1964) Selective suppression of author, title, or the entire citation is supported.

• The Derivative Suit under Japanese Law as a Measure of Corporate Governance (LL.M. thesis, School of Oriental and African Studies, 1994)

• Masahiro Homma, (LL.M. thesis, School of Oriental and African Studies, 1994)

• (invisible)

2Notice how the pinpointed citation gives only as many parallel as are specified in the pinpoint. Compare this with the next example.

References

Macauley, Non-Contractual Relations in Business: A Preliminary Study, 28 Am. Soc. Rev. 55 (1963). Appleyard, Why Major should heed the press, The Independent, p. 21 Jul. 7, 1995. Haley, Japan’s New Land and House Lease Law, in Land Issues in Japan: A Policy Failure? 149 (1992). M. Goosens, F. Mittelbach, and A. Samarin, The L^{A}TEX Companion: The L^{A}TEX Companion (1994). M. Goosens, F. Mittelbach, and A. Samarin, Bibliography Generation, Chapter 13 of The L^{A}TEX Companion 371 (1994). Sir George Bailey Sansom, The First Japanese Constitution: A Lecture (1938). Masahiro Homma, The Derivative Suit under Japanese Law as a Measure of Corporate Governance (LL.M. thesis, School of Oriental and African Studies, 1994). Heap v Motorists’ Advisory Agency Ltd, [1923] 1 KB 577. Bradshaw v United States, 231 Ct.Cl. 144; 683 F.2d 365 (1982). J. M. Maki, Court and Constitution in Japan: Selected Supreme Court Decisions, 1948-1960 (1964).